OpenStudy (allieeslabae):

12. What is the equation of the line perpendicular to the line y = x – 2 that passes through the point (4,3)? y = –4/9 x – 2 y = 9/4 x + 6 y = – 9/4 x + 12 y = – 9/4 x – 12

OpenStudy (anonymous):

Hi

OpenStudy (allieeslabae):

Hey do you know how to solve?

OpenStudy (anonymous):

add me as a fan so I can pm you then I will help you

OpenStudy (anonymous):

To find the the perpendicular line of y = x – 2, first we should start with the slope. The slope of the first equation is x. The slope of an equation perpendicular to the original will have the opposite reciprocal of the original slope. So we take the reciprical of x which is 1/x. The opposite (or negative) of 1/x would be -1/x. Make sense so far?

OpenStudy (allieeslabae):

Okay so slope is x1 x1/ y1 y2 right?

OpenStudy (anonymous):

Could you explain where you got that answer from?

OpenStudy (allieeslabae):

Thats how you find the slope right? Sorry I meant (x1, y1) and (x 2, y 2)

OpenStudy (anonymous):

That is how you would find the slope when given to pairs of coordinates. For example, if you were given the coordinates (3,4) and (1,2) \[\frac{ y2-y1 }{ x2-x1 } \rightarrow \frac{ 4-2 }{ 3-1 } \rightarrow \frac{ 2 }{ 2 } =1\] But to find the slope perpendicular to the original slope, you don't have to use the formula.

OpenStudy (allieeslabae):

Okay im confused wht do i use then?

OpenStudy (anonymous):

It's just what I said earlier "the opposite reciprocal". That's it. :)

OpenStudy (allieeslabae):

Okay thanks you. Ill try that!

OpenStudy (anonymous):

So you know how I got -1/x right?

OpenStudy (allieeslabae):

nope, but I have to go to my tutor, Ill catch her up on what you told me! Thanks though!

OpenStudy (anonymous):

You're welcome. And good luck. :)